Gina has 24 feet of fence.
Largest rectangler area possible?

Answer by checkley77(12844)

The largest area would be a square.
Thus each sid ewould be 24/4=6
6*6=36 ft^2
Proof:
Try any other length & width that add up to 12.
Example 5.9+6.1=12
5.9*6.1=35.99